Andy is twice as old as Kate. In 6 years, their ages will total 60. How old is each now?
This is what I did. However, when I did the math for how old A and K are NOW, it didn't work out.
Let Andy = A
Let Kate = K
A = 2K
2K + K = 60
3K = 60
K = 20
Kate will be 20 and Andy will be 40 in 6 years.
Now they are: Kate = 20 - 6, Andy = 40 - 6
Kate is 14 now and Andy is 34 now.
HOWEVER, if this is true, Andy is NOT twice as old as Kate.
I don't understand.

Answer by Alan3354(30968)

Andy is twice as old as Kate. In 6 years, their ages will total 60. How old is each now?
This is what I did. However, when I did the math for how old A and K are NOW, it didn't work out.
Let Andy = A
Let Kate = K
A = 2K
----------- It's ok to this point.
2K + K = 60 *********** Not now, in 6 years.
========================
A+6 + (K+6) = 60
2K + 6 + K + 6 = 60
3K = 48
K = 16 yrs
A = 32